# Deep learning-based species-area models reveal multi-scale patterns of species richness and turnover

### v1

@@ Abstract
The number of species within ecosystems is influenced not only by their intrinsic characteristics but also by the spatial scale considered. As the sampled area expands, species richness increases, a phenomenon described by the species-area relationship (SAR). The accumulation dynamics of the SAR results from a complex interplay of biotic and abiotic processes operating at various spatial scales. However, the challenge of collecting exhaustive biodiversity records across spatial scales has hindered a comprehensive understanding of these dynamics. Here, we develop a deep learning approach that leverages sampling theory and small-scale ecological surveys to spatially resolve the scale-dependency of species richness. We demonstrate its performance by predicting the species richness of vascular plant communities across Europe, and evaluate the predictions against an independent dataset of plant community inventories. Our model improves species richness estimates by 32% and delivers spatially explicit patterns of species richness and turnover for sampling areas ranging from square meters to hundreds of square kilometers. Explainable AI techniques further disentangle how drivers of species richness operate across spatial scales. The ability of our model to represent the multi-scale nature of biodiversity is essential to deliver robust biodiversity assessments and forecasts under global change.
@@ 1 Introduction
Biodiversity is experiencing unprecedented changes due to global change stressors (IPBES, 2019), with impacts that critically vary across spatial scales (Smart et al., 2006). Agriculture, for example, can increase local diversity through biotic homogenization, but causes biodiversity loss at regional scales by degrading natural habitats (Fahrig, 1997; Maxwell et al., 2016; Simkin et al., 2022; Keil et al., 2015). Similarly, climate change drives divergent richness patterns across spatial scales by altering species distributions and community assembly processes (IPBES, 2019). Assessing how diversity responds to global change across spatial scales is therefore crucial for evaluating the state and trends of biodiversity (Powers and Jetz, 2019; Newbold, 2018; He and Hubbell, 2011). Yet, scaling up small-scale effects from local to regional levels poses a major challenge, resulting in a critical knowledge gap (IPBES et al., 2018).
@@
While the power-law SAR model is a convenient formulation, it relies on the strong assumption of a linear relationship between area and species richness on a log-log scale, which may be considered only valid for a limited range of scenarios and spatial scales (Dengler, 2009; Guilhaumon et al., 2008). In reality, the accumulation of species richness changes in a non-linear fashion as different environmental drivers and ecological barriers are encountered at increasingly larger spatial extents (Conor and McCoy (2013); Borda-de-Água et al. (2012)). As a consequence, the SAR is also location-specific, and depends on biotic and abiotic factors within the area considered (Betts et al. (2014); Matthews et al. (2016); He and Legendre (2002)). Drivers, such as climate, topography, soil conditions, and habitat type shape the spatial distribution of species, affecting the rate at which they accumulate (He and Legendre, 2002; He and Hubbell, 2011; Weigelt and Kreft, 2013; Matthews et al., 2021b; Borda-de-Água et al., 2012; Ulrich et al., 2022; Drakare et al., 2006). Importantly, the effect of these drivers is also scale dependent (Willis and Whittaker, 2002; Turner and Tjørve, 2005; Wiens, 1989).
@@
A number of studies have characterized the effect of environmental drivers on the SAR (Turner and Tjørve, 2005). One branch of research has relied on species inventories from islands or political units to investigate these effects. For instance, Triantis et al. (2003); Kallimanis et al. (2008) showed that incorporating habitat heterogeneity into SAR models better captures spatial variations in species richness, while Kalmar and Currie (2007) extended a power law SAR model by incorporating climate-species interactions to predict global bird species richness. However, the coarse spatial scales of these units limits the characterization of subtle differences in SAR arising at finer spatial grains. Another line of research has focused on developing methods to estimate the SAR from small-scale ecological surveys, some of which are reviewed in (Kunin et al., 2018). These methods are based on sampling theory, which, through explicit or implicit assumptions about the species density distribution, predicts the expected species richness given a certain sampling effort (Elena Schmitz and Rahmann, 2025). Notably, Portier et al. (2022) used small-scale plots that were incrementally aggregated within political units to build country-specific rarefaction curves, predicting expected observed species richness for a given sampling effort. These curves were fitted with a simple statistical model parameterized by environmental features to produce standardized tree-species richness assessments. With a sufficiently expressive statistical model, we could generalize this method to enable the prediction of location and area-specific rarefaction curves. Extrapolating these rarefaction curves to infinite sampling effort could enable the prediction of spatially explicit variations in species richness across spatial scales (Chao et al., 2014; SoberónM. and LlorenteB, 1993; Gotelli and Colwell, 2001).
@@
Here, we present a deep learning approach that leverages small-scale ecological surveys and sampling theory to predict spatially explicit SARs. Specifically, we train a deep neural network to predict the rarefaction curve expected within a spatial unit based on its area and environmental characteristics (see Fig. 1 for an illustration). The neural network parametrizes an asymptotic statistical model that predicts expected species richness for a given sampling effort. By extrapolating this statistical model to its asymptote, we can estimate total species richness within the spatial unit under consideration. We formally justify this approach using sampling theory in Methods. Since our model predicts total species richness within spatial units of arbitrary area, we refer to it as a deep SAR model. We train our deep SAR model using 400k vegetation plots from the European Vegetation Archive, comprising 9,480 distinct species (EVA, Chytrý et al. (2016)), to predict vascular plant species richness across Europe. We evaluate the model’s extrapolation performance using an independent dataset of regional plant inventories from the GIFT database (Weigelt et al., 2020). For environmental features, we calculate the mean and standard deviation of seven bioclimatic variables within each spatial unit, characterizing both mean environmental conditions and environmental heterogeneity. We consider three model variants trained with area only, environmental features only, or both, to systematically evaluate their contributions to predicting species richness. We benchmark our deep SAR model against a traditional non-parametric species richness estimator, and employ Shapley value analysis to disentangle how environmental drivers influence species richness across spatial scales. Finally, we demonstrate the model’s capacity to generate spatially explicit maps of plant species richness and turnover across Europe at multiple spatial scales, revealing the multi-scale aspect of biodiversity patterns that traditional approaches cannot capture.
@@
Figure 1: Deep learning approach for predicting location and area-specific rarefaction curves from small-scale ecological surveys. (a) Data preparation and feature extraction. The top layer shows the location of ecological surveys, which are randomly pooled within spatial units of varying sizes and locations (blue and red boxes) to construct location and area-specific rarefaction curves. The bottom layers show environmental covariates, which we use to derive features characterising the spatial units. (b) Location and area- specific species rarefaction estimation. For each spatial unit, subsampled species counts (scatter points) are fitted with an asymptotic statistical model, which we extrapolate to estimate the total expected species richness under infinite sampling effort. We justify this approach with sampling theory (see Methods for details). A deep neural network learns to predict the parameters of this asymptotic statistical model based on the spatial unit environmental features, enabling the estimation of the rarefaction curve for any location and spatial extent, and the prediction of species richness over increasing area. In practice, we only generate one subsampled species count per spatial unit, but many more spatial units, to span large extents efficiently.
@@ 2.1 Model performance
The deep SAR model demonstrated superior performance in interpolating across location and area-specific species rarefaction curves when utilizing environmental features (Fig. 2a). To evaluate the models’ performance, we used held-out vegetation plots (see Biodiversity data. for details on the training and test datasets), assessing their capacity to predict species rarefaction curves for low to moderate sampling effort. We trained the deep SAR using three sets of features: area only, environmental features only, and both area and environmental features. Following Kunin et al. (2018), we employed the root mean squared error (RMSE) to evaluate model performance, measured across five independent training runs. Providing the deep SAR model with statistics on environmental features reduced the RMSE on the EVA test dataset by 43% compared to using area only. The performance of the model informed with both area and environmental features was not significantly improved compared to the model trained with environmental features only. This can be attributed to the high correlation between the standard deviations of environmental features and the spatial unit area (see correlation matrix in Fig. S7). The best models showed a median relative bias of -19% in predicting the species rarefaction curves on the EVA-based test dataset (Fig. 2b), indicating a moderate tendency to overpredict species richness. This bias was more pronounced for low values of species richness and decreased as species richness increased.
@@
The model trained with both area and environmental features together demonstrated the best performance in extrapolating to predict total species richness. Total species richness was obtained by extrapolating the predicted rarefaction curve towards its asymptote, assuming infinite sampling effort. We matched these predictions against species richness derived from species inventories obtained from the GIFT database. Overall, the best extrapolating model was the one leveraging both area and environmental features. The RMSE of this model was 23% lower than that of the model trained with area only. Contrary to its interpolation performance, we found a significant advantage in using area together with environmental features over environmental features only (performance gain of 12%, $p$-value $<0.001$ under t-test). The improved performance provided by the sampling unit area can be interpreted as an effect of "area per se" (Connor and McCoy, 1979). Along this mechanism, larger areas promote higher species richness, even when discounting for increased diversity in environmental niches, as they increase the population size, allowing more individuals to be sampled, and as they bridge barriers to dispersal (Drakare et al., 2006). We note that the distribution of data points in the GIFT dataset is biased towards larger areas (in the order of magnitude of a country’s area, see Fig. S6). At this scale, the summary statistics of environmental features used to inform the model are less informative than for smaller areas (see Contribution of area and environment across spatial scales and Fig. 3) and where species richness shows less variation. We expect that more fine-grained data points or the use of higher-order summary statistics for the environmental features would give a larger advantage to the deep SAR model trained with environmental features.
Last but not least, the deep SAR model significantly outperformed the Chao2 estimator on extrapolation performance (gain of 32%, $p$-value $<0.001$ under t-test) and provided comparably less biased estimates of species richness (median relative bias of 22% compared to 52% for the Chao2 estimator).
@@
Figure 2: Ablation study and performance evaluation of the deep SAR model. (a) Interpolation performance of the deep SAR model for low to moderate sampling effort, as measured by the root mean squared error (RMSE), trained with each set of features, on the held-out plot vegetation test dataset. Different lower letters above the boxplots indicate that performance are significantly different from each other (Posthoc Tukey HSD test). (b) Observed species richness against values predicted by the best model (area + environment). (c) Extrapolation performance of the deep SAR model, trained with each set of features, and compared against the non-parametric Chao2 estimator, on the GIFT-based test dataset. (d) Observed species richness against values predicted by the best model (area + environment).
@@
In summary, these results suggest that (i) our model effectively predicts the rarefaction curves associated with spatial units of arbitrary location and extent, providing the expected species richness given a certain sampling effort, and (ii) reliably extrapolates to predict total species richness. Hence, the proposed approach successfully accounts for variations in SAR due to environmental drivers, capturing species richness patterns across scales.
@@ 2.2 Contribution of area and environment across spatial scales
Disentangling the effect of spatial unit area on species richness from environmentally-related confounding factors has been a challenging task in macroecology, as these factors act differently depending on spatial scale. To quantify the relative contributions of environmental features versus area, we calculated the relative Shapley values (Lundberg and Lee, 2017) for each predictor class (area of the spatial unit considered, mean and heterogeneity of environmental conditions), stratified by sampling area. Specifically, we calculated absolute Shapley values for each predictor class and normalized them so that the sum of Shapley values for all features equals one for each prediction of species richness. This approach allows us to characterize the relative importance of each predictor class across varying sampling areas, as shown in Fig. 3.
@@
Figure 3: Relative importance of area and environmental conditions in predicting species richness. This plot shows the relative importance as indicated by Shapley values for the spatial unit area and environmental features, grouped into two classes: mean environmental conditions and environmental heterogeneity. Higher values indicate a greater contribution of the respective predictor group to predicting species richness at any given spatial grain. Plain lines indicate the mean relative absolute Shapley values across all predictions, while shaded areas represent the standard deviation.
@@
Mean environmental conditions were the dominant drivers of species richness across all spatial scales investigated. This finding supports the environmental filtering hypothesis, wherein mean environmental conditions constrain species richness by limiting species establishment based on physiological tolerances (Whittaker, 1972; Rahbek, 2005). As the spatial unit area increases, the relative contributions of both area and environmental heterogeneity increased, while the influence of mean environmental conditions decreased. Environmental heterogeneity promotes higher species richness by increasing niche diversity, which supports greater species turnover (Stein et al., 2014; Qian and Ricklefs, 2012; Turner and Tjørve, 2005). Meanwhile, larger sampling areas can enhance species richness by allowing more individuals to be sampled ("area per se" effect, see previous section). We note that habitat area is strongly correlated with the standard deviation of bioclimate variables in our dataset (see Fig. S7), which may confound the interpretation of area effects versus environmental heterogeneity. These findings quantify the scale-dependent nature of biodiversity drivers (Rahbek, 2005) and provide quantitative evidence for the mechanisms underlying species richness patterns across spatial scales (Field et al., 2009).

### v2

# Multi-scale species richness estimation with deep learning

@@ Abstract
Biodiversity assessments are critically affected by the spatial scale at which species richness is measured. How species richness accumulates with sampling area depends on natural and anthropogenic processes whose effects can change depending on the spatial scale considered. These accumulation dynamics, described by the species-area relationship (SAR), are challenging to assess because most biodiversity surveys are restricted to sampling areas much smaller than the scales at which these processes operate. Here, we combine sampling theory and deep learning to predict local species richness within arbitrarily large sampling areas, enabling for the first time to estimate spatial differences in SARs. We demonstrate our approach by predicting vascular plant species richness across Europe and evaluate predictions against an independent dataset of plant community inventories. The resulting model, named deep SAR, delivers multi-scale species richness maps, improving coarse grain richness estimates by 32% compared to conventional methods, while delivering finer grain estimates. Additional to its predictive capabilities, we show how our deep SAR model can provide fundamental insights on the multi-scale effects of key biodiversity processes. The capacity of our approach to deliver comprehensive species richness estimates across the full spectrum of ecologically relevant scales is essential for robust biodiversity assessments and forecasts under global change.
Significance statement. Understanding how biodiversity responds to global change requires assessing species richness across spatial scales. Yet, most biodiversity data are collected at local scales, while the ecological processes that shape richness often act regionally or globally. This mismatch makes it difficult to track biodiversity change effectively. Here, we bridge this gap by integrating sampling theory with deep learning to predict species richness across scales. Using $\sim$490k vegetation surveys, we map vascular plant richness across Europe and validate our predictions with independent data. Our model, named deep SAR, improves large-scale richness estimates by 32% compared to standard methods, while revealing the multi-scale nature of biodiversity patterns. This work delivers a scalable framework for monitoring and forecasting biodiversity under global change.
@@ 1 Introduction
Biodiversity is experiencing unprecedented changes due to global change stressors (IPBES, 2019), with impacts that critically vary across spatial scales (Smart et al., 2006). Agriculture, for example, can increase local diversity through biotic homogenization, but causes biodiversity loss at regional scales by degrading natural habitats (Fahrig, 1997; Maxwell et al., 2016; Simkin et al., 2022; Keil et al., 2015). Similarly, climate change drives divergent richness patterns across spatial scales by altering species distributions and community assembly processes (IPBES, 2019). Assessing how diversity responds to global change across spatial scales is therefore crucial for evaluating the state and trends of biodiversity (Powers and Jetz, 2019; Newbold, 2018; He and Hubbell, 2011; Rahbek, 2005). Yet, scaling up small-scale effects from local to regional levels poses a major challenge, resulting in a critical knowledge gap (IPBES et al., 2018).
@@
While the power-law SAR model is a convenient formulation, it relies on the strong assumption of a linear relationship between area and species richness on a log-log scale, which may be considered only valid for a limited range of scenarios and spatial scales (Dengler, 2009; Guilhaumon et al., 2008). In reality, the accumulation of species richness changes in a non-linear fashion as different environmental drivers and ecological barriers are encountered at increasingly larger spatial extents (Conor and McCoy (2013); Chisholm et al. (2025); Borda-de-Água et al. (2025)). As a consequence, the SAR is also location-specific, and depends on biotic and abiotic factors within the area considered (Betts et al. (2014); Matthews et al. (2016); He and Legendre (2002)). Drivers, such as climate, topography, soil conditions, and habitat type shape the spatial distribution of species, affecting the rate at which they accumulate (He and Legendre, 2002; He and Hubbell, 2011; Weigelt and Kreft, 2013; Matthews et al., 2021b; Chisholm et al., 2025; Borda-de-Água et al., 2025; Ulrich et al., 2022; Drakare et al., 2006). Importantly, the effect of these drivers is also scale dependent (Willis and Whittaker, 2002; Turner and Tjørve, 2005; Wiens, 1989).
@@
A number of studies have characterized the effect of environmental drivers on the SAR (Turner and Tjørve, 2005). One branch of research has relied on exhaustive species inventories from islands or political units to investigate these effects. For instance, Triantis et al. (2003); Kallimanis et al. (2008) showed that incorporating habitat heterogeneity into SAR models better captures spatial variations in species richness, while Kalmar and Currie (2007) extended a power law SAR model by incorporating climate-species interactions to predict global bird species richness. However, the coarse spatial scales of these spatial units limits the characterization of subtle differences in SAR arising at finer spatial grains.
To resolve species richness at finer spatial scales, another line of research has relied on upscaling methods to estimate species richness from partially sampled spatial units (Kunin et al., 2018; Portier et al., 2022; kusumoto2023). These approaches use sampling theory, which, by making explicit or implicit assumptions about the species density distribution, enable the estimation of the total species richness within the spatial unit considered under asymptotic sampling effort, based on a limited number of small-scale ecological surveys (Chao et al., 2014; Elena Schmitz and Rahmann, 2025). One prominent method involves fitting a parametric function to the species rarefaction curve, which describes how species richness increases with sampling effort within a given spatial unit (see Fig.˜1 for an illustration). Unlike the SAR, which typically relates richness to a contiguous area, the rarefaction curve describes how expected richness grows with sampling effort, where samples are not necessarily contiguous. The rarefaction curve is bounded by total species richness within the spatial unit, which can be estimated by extrapolation under asymptotic sampling effort (Soberón M. and Llorente B., 1993; Gotelli and Colwell, 2001). Conditioning this parametric function on the spatial unit’s area and associated environmental covariates could theoretically enable the prediction of location- and area-specific rarefaction curves; by extrapolation, it could provide location-specific estimates of species richness for arbitrary spatial grains.
@@
Here, we present a deep learning approach that leverages small-scale ecological surveys and sampling theory to predict spatially explicit SARs. We train a deep neural network to predict the species rarefaction curve expected within a spatial unit based on its area and associated environmental covariates (see Fig.˜1 for an illustration). The neural network parametrizes a parametric function that predicts expected species richness for a given sampling effort. By extrapolating this statistical model to its asymptote, we can estimate total species richness within the spatial unit. We formally justify this approach using sampling theory in Methods. Since our model can predict total species richness within spatial units of arbitrary area, we refer to it as a deep SAR model. Using a large dataset of $\sim$490k vegetation plots from the European Vegetation Archive comprising 9,480 distinct species (EVA, Chytrý et al. (2016)), we show how our deep SAR model enables multi-scale estimates of vascular plant species richness across Europe. We evaluate the model’s interpolation performance using a held-out dataset, and its extrapolation performance using an independent dataset of regional plant inventories from the GIFT database (Weigelt et al., 2020). For environmental features, we calculate the mean and standard deviation of seven bioclimatic variables within each spatial unit, characterizing respectively mean environmental conditions and environmental heterogeneity. We consider three model variants trained using area only, environmental features only, or both, to systematically evaluate their contributions to predicting species richness. We benchmark our deep SAR model against a conventional non-parametric species richness estimator, and employ Shapley value analysis to disentangle how environmental drivers influence species richness across spatial scales. Finally, we demonstrate the model’s capacity to generate spatially explicit maps of plant species richness and turnover across Europe at multiple spatial scales, revealing the multi-scale aspect of biodiversity patterns.
@@
Figure 1: Deep learning approach for predicting location and area-specific rarefaction curves from small-scale ecological surveys. (a) Data preparation and feature extraction. The green cells on the top layer correspond to the location of ecological surveys, which are randomly pooled within spatial units of varying sizes and locations (blue and red boxes) to construct location and area-specific rarefaction curves. The bottom layers show environmental covariates, which we use to derive features characterising the spatial units. (b) Location and area- specific species rarefaction estimation. For each spatial unit, subsampled species counts (scatter points) are fitted with an asymptotic statistical model, which we extrapolate to estimate the total expected species richness under infinite sampling effort. We justify this approach with sampling theory (see Methods for details). A deep neural network learns to predict the parameters of this asymptotic statistical model based on the spatial unit environmental features, enabling the estimation of the rarefaction curve for any location and spatial extent, and the prediction of species richness over increasing area. In practice, we only generate one subsampled species count per spatial unit, but many more spatial units, to span large extents efficiently.
@@ 2.1 Model performance
Environmental features could significantly improve the estimation of species rarefaction curves (Fig.˜2a). We trained three model variants using different sets of features: area only, environmental features only, and both area and environmental features. Using held-out vegetation plots (see Biodiversity data. for details on the training and test datasets), we assessed the ability of the models to predict species rarefaction curves across spatial units of varying location and area, for low to moderate sampling effort. Following Kunin et al. (2018), we calculated the root mean squared error (RMSE) of the predictions, measured across five independent training runs. Providing the models with statistics on environmental covariates reduced the RMSE on the EVA test dataset by 43% compared to using area only. The performance of the model informed with both area and environmental features was not significantly improved compared to the model trained with environmental features only. This can be attributed to the high correlation between the standard deviations of environmental features and the spatial unit area (see correlation matrix in Fig.˜S3). The best models showed a median relative bias of 19% in predicting the species rarefaction curves on the EVA-based test dataset (Fig.˜2b), indicating a moderate tendency to overpredict species richness. This bias was more pronounced for low values of species richness and decreased towards 0 as species richness increased (see Fig.˜S4). This result demonstrates the capability of our approach to resolve fine-grained variations in species rarefaction curves.
@@
The model variant trained with both area and environmental features together demonstrated superior performance in extrapolating to predict total species richness (Fig.˜2a). Total species richness was obtained by extrapolating the predicted rarefaction curve towards its asymptote, assuming infinite sampling effort. We matched these predictions against species richness derived from species inventories obtained from the GIFT database. The RMSE of the model trained with area and environmental features was 23% lower than that of the model trained with area only. Contrary to its interpolation performance on the EVA dataset, we found a significant advantage in using area together with environmental features over environmental features only (performance gain of 12%, $p$-value <0.001<0.001 under t-test). The improved performance provided by the sampling unit area can be interpreted as an effect of "area-per se" (Connor and McCoy, 1979). Along this mechanism, larger areas promote higher species richness independently of environmental niche diversity (Drakare et al., 2006), where increasing area alone allows for more individuals to be sampled and bridging dispersal barriers. We note that the distribution of data points in the GIFT dataset is biased towards larger areas (in the order of magnitude of a country’s area, see Fig.˜S2). At this scale, the summary statistics of environmental features used to inform the model are less informative than for smaller areas (see Contribution of area and environment across spatial scales and Fig.˜3), and species richness shows less variation. We expect that more fine-grained species inventories or the use of higher-order summary statistics for the environmental features would give a larger advantage to the model trained with environmental features. It is worth noting that the model significantly outperformed the Chao2 estimator in extrapolation performance, achieving a 32% reduction in RMSE ($p$-value <0.001<0.001, t-test) and providing less biased estimates of species richness (median relative bias of -22% compared to -52% for the Chao2 estimator).
@@
Figure 2: Ablation study and performance evaluation of the deep SAR model. (a) Interpolation performance of the model for low to moderate sampling effort, as measured by the root mean squared error (RMSE), trained with each set of features, on the held-out plot vegetation test dataset (EVA test dataset). Different lower letters above the boxplots indicate that performance are significantly different from each other (Posthoc Tukey HSD test). (b) Observed species richness against values predicted by the best model (area + environment). (c) Extrapolation performance of the model, trained with each set of features, and compared against the non-parametric Chao2 estimator, on the species inventories test dataset (GIFT test dataset). (d) Observed species richness against values predicted by the best model (area + environment).
@@
In summary, these results suggest that (i) our model effectively predicts the rarefaction curves associated with spatial units of arbitrary location and extent, providing the expected species richness given a certain sampling effort, and (ii) reliably extrapolates to predict total species richness. As a consequence, the model can predict variations in species richness across locations and spatial scales, capturing location-specific species-area relationships. In the following, we refer to the model trained with area and environmental features as the deep SAR model.
@@ 2.2 Contribution of area and environment across spatial scales
Disentangling the effect of spatial unit area on species richness from environmentally-related confounding factors has been a challenging task in macroecology, as these factors act differently depending on spatial scale. To quantify the relative contributions of environmental features versus area, we calculated the Shapley values associated with the prediction of total species richness on the EVA test dataset for each predictor class (area of the spatial unit considered, mean and heterogeneity of environmental conditions), stratified by sampling area. Shapley values are commonly used in explainable artificial intelligence to measure the contribution of each feature to the model’s output by evaluating all possible feature combinations (Lundberg and Lee, 2017). Specifically, we calculated absolute Shapley values for each predictor class and normalized them so that the sum of Shapley values for all features equals one for each species richness estimates. This approach allowed us to characterize the relative importance of each predictor class across varying sampling areas, as shown in Fig.˜3.
@@
Figure 3: Relative importance of area and environmental features in predicting total species richness. Shapley values indicate the contribution of each predictor class, i.e., spatial unit area, mean environmental conditions (mean of environmental covariates), and environmental heterogeneity (standard deviation of environmental covariates), to total species richness estimates at different spatial grains. Lines show mean relative absolute Shapley values; shaded areas show standard deviation.
@@
Mean environmental conditions were contributing the most to species richness predictions across all spatial scales investigated. This finding supports the environmental filtering hypothesis, wherein mean environmental conditions constrain species richness by limiting species establishment based on physiological tolerances (Whittaker, 1972; Vellend, 2010). As the spatial grain increased, the relative contributions of both area and environmental heterogeneity increased, while the contribution of mean environmental conditions decreased. The contribution of spatial unit area was particularly pronounced, with a sustained increase as spatial grain increased, whereas the contribution was overall least pronounced (see Biodiversity data.). This finding resonates again with the "area-per se" effect (see previous section), where larger areas can support more species simply by providing more space and resources, independent of environmental conditions. Meanwhile, the increasing contribution of environmental heterogeneity at larger spatial grains aligns with the "habitat-diversity hypothesis", wherein environmental heterogeneity promotes higher species richness by increasing niche diversity, which supports greater species turnover (Stein et al., 2014; Qian and Ricklefs, 2012; Turner and Tjørve, 2005). We note that environmental heterogeneity naturally began to rise only at spatial grains larger than the environmental covariate resolution; using environmental covariates with a finer resolution would likely change this pattern. Overall, these findings illustrate the scale-dependent nature of biodiversity drivers (Rahbek, 2005) and help quantify the relative contributions of mechanisms underlying species richness patterns across spatial scales (Field et al., 2009).